OpenStudy (anonymous):

can someone help with my cal 3? i posted a question and im still stuck.

OpenStudy (anonymous):

yeah whats up

OpenStudy (anonymous):

ok thanks im going to type it now...

OpenStudy (anonymous):

\[\int\limits_{0}^{6}\int\limits_{?}^{?}\]

OpenStudy (anonymous):

\[\int\limits_{0}^{5}\int\limits_{0}^{y}(siny/y)\]

OpenStudy (anonymous):

dx dy ?

OpenStudy (anonymous):

dxdy

OpenStudy (anonymous):

with respect to dx, the integral of sin y/y is just a constant so you get [ sin y/y *x ] and you integrate from y to 0

OpenStudy (anonymous):

ok so it would be integrate from 5 to 0 (siny^2/2)?

OpenStudy (anonymous):

no you didnt plug in y to 0 for x

OpenStudy (anonymous):

integral [ sin y/y [ y - 0] dy

OpenStudy (anonymous):

when you plug y in for x will the bottom y cancel out?

OpenStudy (anonymous):

\[\int\limits_{0}^{y}\]

OpenStudy (anonymous):

im lost

OpenStudy (anonymous):

its not hard

OpenStudy (anonymous):

this latex thing is killing me

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

it can you please work it out for me?

OpenStudy (anonymous):

\[\int\limits_{0}^{5}\int\limits_{y}^{0} siny/y dx dy = \]

OpenStudy (anonymous):

forget it, im sick of this , find someone else. this latex is retarded

OpenStudy (anonymous):

since siny/y is a constant wouldn't you multiply it times x and then plug y in for x?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

go ahead and do that

OpenStudy (anonymous):

\[\int\limits_{0}^{5}\]\[\int\limits_{0}^{y}\]

OpenStudy (anonymous):

\[\int\limits_{?}^{?}\int\limits_{?}^{?}\]

OpenStudy (anonymous):

ok after i did that i got \[\int\limits_{0}^{5}sinydy \]

OpenStudy (anonymous):

anyone?

OpenStudy (anonymous):

this is baby calculus, cmon